the length of a garden is 10 m more than twice the width. The are is 120m2. What are the dimensions of the garden to the nearest tenth metre?
Answer by Boreal(15235)

W=x
L=2x+10
area is 120 m^2
Therefore, (2x+10)(x)=120
2x^2+10x=120
2x^2+10x-120=0
divide by 2
x^2+5x-60=0
x=(1/2) { -5 +/- sqrt (25+240)}, and use positive root only
sqrt 265=16.279
x=(1/2)(-5+16.279)=(1/2)(11.279)=5.64 m
2x+10=21.28
convert to nearest tenth of a meter
5.6*21.3=119.28
rounded to two decimal places it is 120.02.
The dimensions are 5.6 m x 21.3 m